Given log 2 = 0.3010 and log 3 = 0.4771
What is log 1/25

Answer by lwsshak3(11628)

Given log 2 = 0.3010 and log 3 = 0.4771
What is log 1/25
log 1/25=log 4/100= 2log2-log100=2*.3010-2=.6020-2=-1.398